OpenStudy (anonymous):

Find the volume of the sphere to the nearest whole number. Use p = 3.14. A. 131 in.^3 B. 393 in.^3 C. 4,187 in.^3 D. 523 in.^3

OpenStudy (anonymous):



OpenStudy (anonymous):

Volume = 4/3πr^3 4 4 = ---- * 3.14 * (5)^3 = ----- * 3.14 * 125 = 523.33 3 3 so answer is D. 523 in.^3